# Beyond-$\Lambda$CDM Cosmologies with Bayesian Neural Networks II: Massive Neutrinos, Baryonic Feedback, and the Theoretical Error

### v1

@@ 4.5 Scale cuts
4.6 Priors on Cosmological Parameters
To test the effect of priors on the classification we train an EE2 network with $\Omega_{\rm b}$ values from a very narrow distribution with a standard deviation of 0.001 instead of 0.016. The results are shown in Figure 8. If we test this model with data based on the wider $\Omega_{\rm b}$ prior then the total test accuracy gets reduced by 11%. This is a indication that the network cannot extrapolate far beyond the learned parameter range. Many spectra get instead classified as random because the random spectra are based on $\Lambda$CDM spectra with the full $\Omega_{\rm b}$ parameter range. For very strong deviations from Planck’s best fit parameters it might be justified to flag them as new physics in the random class that is representing other unconsidered models. However in general, we can conclude that the influence of cosmological parameter values outside of the prior of the training data have a significant influence on the classification. Such a behaviour is not surprising and was expected both from a machine learning and cosmological perspective. This test simply warns that the prior for the training data has to be selected with care in order to maximise our confidence in the network’s predictions.
@@ 5 Conclusions
Forthcoming Stage-IV cosmological surveys will perform measurements with unprecedented accuracies, in particular at nonlinear scales. We have improved a novel machine learning approach to detect beyond-standard-model physics in the data using a Bayesian Neural Network. The purpose of this machine-learning based method is not the replacement of an MCMC analysis, but the development of a tool that can reduce the vast model space of $\Lambda$CDM modifications with high time efficiency. This includes signs for the preference of $\Lambda$CDM or hints at new physics beyond the considered cosmological classes. Once preferred models are identified, an MCMC analysis can be used to obtain parameter constraints from the selected subset. This approach reduces the data analysis time in the search for new physics without the need for biased assumptions on the model or parameter space.
@@
Our Bayesian Cosmological Network (BaCoN-II) contains convolutional layers that perform an image analysis of the 2-dimensional matter power spectrum. It uses the information in the shape of the power spectrum to classify it into one of five classes: $\Lambda$CDM, $f(R)$, $w$CDM, DGP gravity and a ‘random’ class accounting for unknown physics. Our baseline model is able to distinguish between these classes without any major degeneracies and has a total classification accuracy of $\sim$ 95%.
@@
There are some limitations to the ability of the network to generalise: It is dependent on the choices for the parameter ranges used in the training data and has only a limited ability to extrapolate beyond these priors. This effect was to be expected and might be of use to flag strong deviations from Planck’s best fit parameters. However in general one has to be aware that the selection of priors matters. Furthermore, the influence of baryonic effects, massive neutrinos and power spectrum-specific characteristics is stronger than features generated by modified gravity and dynamic dark energy theories. This can lead to a wrong classification as ‘random’ if one of the effects is missing in the training data. Hence, a consistent construction of the random class based on accurate matter power spectrum is important. However, as the random class represents other feasible models beyond the selected ones this classification shows that BaCoN-II can successfully flag the detection of new physics.

### v2

@@ 4.3 Massive Neutrinos and Baryonic Feedback
We have done a more specific test to understand how the network is able to distinguish between $f(R)$ and massive neutrinos, given that both of these can have strong scale-dependent effects. For that, we’ve trained a two-class model only with $\Lambda$CDM and $f(R)$ data to see if the effect of massive neutrinos in a $\Lambda$CDM cosmology can get confused with $f(R)$. Figure 12 shows two confusion matrices produced using two distinctly trained models. Both models have been trained and tested with $f(R)$ spectra without massive neutrinos but the $\Lambda$CDM in training varies. It either included or didn’t include massive neutrinos. The test is then performed with massive neutrinos in the $\Lambda$CDM testing data. If we don’t account for massive neutrinos in training, their effect gets mistaken as $f(R)$. But if we train with neutrinos, there is near to no degeneracy. This means that they have a broadly similar effect on the shape of the power spectrum, so they can get mistaken for one another in the lack of more class-specific information. However, when trained with spectra including both effects, then the specific imprints of $f(R)$ and massive neutrinos in the power spectrum are distinctive enough to confidently separate them.
Figure 19 shows that the changes that $f(R)$ and massive neutrinos induce in the power spectrum have different shapes on nonlinear scales. So the classification ability of the network most likely originates from including these highly nonlinear scales in the training data. To understand more about the specific patterns that the network is looking for, we would need methods that visualise the decision making process of the network, like Class activation maps. We will implement explainable AI methods like this in our next work. Additionally, this test emphasises how important it is to accurately model all of the physics on nonlinear scales as this might otherwise lead to a false detection of modified gravity, for example.
Figure 12: Influence of massive neutrinos: confusion matrices of two-class models trained on $f(R)$ spectra without massive neutrinos and $\Lambda$CDM data with (right) or without (left) massive neutrinos (with a 5% theoretical error component). The test data is the same in both cases, it contains $f(R)$ without and $\Lambda$CDM data with massive neutrinos.
@@ 4.5 Scale Cuts
4.6 Prior effects and Out-Of-Distribution samples
To test the effect of priors on the classification we train an EE2 network with $\Omega_{\rm b}$ values from a very narrow distribution with a standard deviation of 0.001 instead of 0.016. The results are shown in Figure 11. If we test this model with data based on the wider $\Omega_{\rm b}$ prior then the total test accuracy gets reduced by 11%. This is a indication that the network cannot extrapolate far beyond the learned parameter range. Many spectra get instead classified as random because the random spectra are based on $\Lambda$CDM spectra with the full $\Omega_{\rm b}$ parameter range. For very strong deviations from Planck’s best fit parameters it might be justified to flag them as new physics in the random class that is representing other unconsidered models. However in general, we can conclude that the influence of cosmological parameter values outside of the prior of the training data have a significant influence on the classification.
Figure 17: Classification of Out-Of-Distribution samples: The histograms show the distribution of values for cosmological parameters in the $\Lambda$CDM data of the baseline EE2 training data (dark gray) and test data (gray) as well as a set with Out-Of-Distribution (OOD) $\Lambda$CDM data (orange). The bar height represents the full count for test and OOD data, and has been scaled for the 20,000 training samples. The second row shows the $\Lambda$CDM-classification accuracies when the baseline EE2-based model is tested with $\Lambda$CDM spectra that are generated from these OOD parameters. The last row features the $\Lambda$CDM-classification accuracies of the In-Distribution baseline test data as shown in Figure 8.
To further probe the extent of this effect for our baseline model we have run a number of tests with Out-Of-Distribution (OOD) data. For that we have produced 200 outliers of the parameter distributions in $\Omega_{\rm m}$, $H_{0}$, $n_{s}$ and $A_{s}$. They are sampled to be $3\sigma$ outliers from the $\Lambda$CDM training set used for our baseline EE2 model. When computing the $\Lambda$CDM power spectra for these samples we only use the OOD value for one of the parameters at a time while sampling the other parameters as normal.
The parameter distributions used in the training data and testing data of the baseline model are shown in Figure 17. We rescaled the height of the bars for the training data to account for the larger number of training samples. The parameter values of the OOD test set are shown in orange. The corresponding confusion matrix for the $\Lambda$CDM class is shown below the respective histogram. Even though all the $\Lambda$CDM spectra are produced from $3\sigma$ outliers, the $\Lambda$CDM true-positive accuracy does not drop by more than 15% compared to the standard test data (see Figure 8). This is a high performance given that all test samples have one parameter that is a significant outlier of the training parameter distribution. It shows that the trained network is fairly robust to these type of OOD samples and our chosen parameter priors.
It should be noted, that none of the OOD samples is classified as random. Hence, the network is able to distinguish between cosmological parameter induced patterns in the spectrum and completely different behaviours expected from a different class of cosmological model. However, we can see that existing degeneracies are worsened. Especially for the $n_{s}$ and $A_{s}$ outliers we see an increased mis-classification of $\Lambda$CDM as $w$CDM. $w$CDM affects the amplitude of the spectrum as well as its scale-dependency and so this degeneracy is expected.
Despite the relative robustness to OOD samples, this exercise shows that priors should be chosen carefully and a wider choice is advantageous. Such a behaviour is not surprising and was expected both from a machine learning and cosmological perspective. In our case, we have chosen Planck and Stage-IV galaxy survey based priors for the training data that assume LSS and the CMB are consistent. This approach should encompass the majority of currently considered cosmological models.
@@ 5 Conclusions
Forthcoming Stage-IV cosmological surveys will perform measurements with unprecedented accuracies, in particular at nonlinear scales. We have improved a novel machine learning approach to detect beyond-standard-model physics in the data using a Bayesian Neural Network. The purpose of this machine-learning based method is not the replacement of a MCMC analysis, but the development of a tool that can reduce the vast model space of $\Lambda$CDM modifications with high time efficiency. This includes signs for the preference of $\Lambda$CDM or hints at new physics beyond the considered cosmological classes. Once preferred models are identified, a MCMC analysis can be used to obtain parameter constraints from the selected subset. This approach reduces the data analysis time in the search for new physics without the need for biased assumptions on the model or parameter space.
@@
Our Bayesian Cosmological Network (BaCoN-II) contains convolutional layers that perform an image analysis of the matter power spectrum. It uses the information in the shape of the power spectrum to classify it into one of five classes: $\Lambda$CDM, $f(R)$, $w$CDM, DGP gravity and a ‘random’ class accounting for unknown physics. Our baseline model is able to distinguish between these classes without any major degeneracies and has a total classification accuracy of $\sim$ 95%. It is well calibrated with a Static Calibration Error of 1.1%. We have also shown that we can use the distribution of output probabilities of the BNN to estimate the uncertainty of the classification result. This is an advantage over deterministic networks for the analysis of observable data, as we will be able to calculate the confidence for the test result of a single spectrum.
@@
There are some limitations to the ability of the network to generalise: It is able to extrapolate beyond the priors of the training data to a certain extent but is still dependent on the choices for the parameter ranges. This effect was to be expected and might be of use to flag strong deviations from Planck’s best fit parameters. However in general one has to be aware that the selection of priors matters. Furthermore, the influence of baryonic effects, massive neutrinos and power spectrum-specific characteristics is stronger than features generated by modified gravity and dynamic dark energy theories. This can lead to a wrong classification as ‘random’ if one of the effects is missing in the training data. Hence, a consistent construction of the random class based on accurate matter power spectrum is important. However, as the random class represents other feasible models beyond the selected ones this classification shows that BaCoN-II can successfully flag the detection of new physics.